OpenStudy (anonymous):

x^3y'' + x^2y' -4xy = 1 (x>0) variation of parameter again @Outkast3r09

OpenStudy (anonymous):

\[x^3y'' + x^2y' -4xy = 1\]

OpenStudy (anonymous):

@Outkast3r09 like i said huh ? :)

OpenStudy (anonymous):

this one is easier

OpenStudy (anonymous):

divide by x^3

OpenStudy (anonymous):

realy :D

OpenStudy (anonymous):

so it is like previous question

OpenStudy (anonymous):

i could not do it

OpenStudy (anonymous):

you need to figure out one solution to this equation

OpenStudy (anonymous):

and that means

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I figured ito ut your teacher is skipping around the book lol

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

he probably wants you to use cauchy-eulers equation to solve for the roots

OpenStudy (anonymous):

my book doesn't teach cauchy euler until after

OpenStudy (anonymous):

it is also about the variation of parameter

OpenStudy (anonymous):

at least my book says so

OpenStudy (anonymous):

obtain a general solution using the method of the variation of parameter :/

OpenStudy (anonymous):

yes however you need a way to find your solutions... and the only way is to using auxilary equations however in this case auxilary equations will not work

OpenStudy (anonymous):

so you have to use cauchy-eulers method of finding the roots

OpenStudy (anonymous):

In my book it says right here "Also, we can solve the nonhomogeneous equation ax^2y''+bxy'+cy=g(x) by variation of parameters once we have determined the yC

OpenStudy (anonymous):

to use eulers you must have x to the exponent n derivative n

OpenStudy (anonymous):

now i haven't taken th test on this so i'm not 100% how it works you'll have to find a derivation of this however if you let \[y=x^m\] \[y'=mx^{m-1}\] \[y''=(m-1)(m)x^{m-2}\]

OpenStudy (anonymous):

first you need the powers to line up so divide by x to get \[x^2y''+xy'-4y=x^{-1}\]

OpenStudy (anonymous):

then substitute in \[x^2(m(m-1)x^{m-1}+x(m)x^{m-1}-4x^m=0\] keep it homogeneous for complementary

OpenStudy (anonymous):

and the first is supposed to be x^{m-2}

OpenStudy (anonymous):

\[x^{m-2}=x^mx^{-2}\] the x's cancel

OpenStudy (anonymous):

you're left with \[x^m(m^2-m+m-4)\] =\[x^m(m^2-4)\]

OpenStudy (anonymous):

\[m^2-4=0\] m=2 k=1 , m=-2, k =1

OpenStudy (anonymous):

since these are rational different roots you get \[y_c=c_1x^{-2}+c_2x^2\]

OpenStudy (anonymous):

now \[y_1=x^{-2},y_2=x^2\]

OpenStudy (anonymous):

\[y_p=u_1y_1+u_2y_2\]

OpenStudy (anonymous):

make sense so far?

OpenStudy (anonymous):

i am trying :D

OpenStudy (anonymous):

i think that's enough for today tomorrow i will try again :) Good Night in Turkey

OpenStudy (anonymous):

yep if you want to finish this yourself. all you have to do is take your y1 and y2 and computer the W(y1,y2) and then W1(f(x),y2) W2(y1,f(x)) and then use kramers rule u1'=W1/W

OpenStudy (anonymous):